OpenStudy (anonymous):

Find the derivative of the following function: y=(2x+1)^-3 (3x-7)^6

OpenStudy (anonymous):

\[y= (2x + 1)^{-3} (3x-7)^{6}\]

OpenStudy (anonymous):

I am also looking for a formula or something that could explain this derivative easily.

OpenStudy (anonymous):

Product rule, power rule, chain rule.

OpenStudy (anonymous):

Or quotient rule, if you're into that kind of thing

OpenStudy (anonymous):

See, my teacher never thought us the rules, he just kinda.. did them haha

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[ \large \text{Differentiation Rules} \]Let \(c\) be constant with respect to \(x\). Let \(u\) and \(v\) vary with respect to \(x\). \[ \begin{array}{r|ccc|l} 1&d(c) &=& 0&\text{Constants}\\ \ 2&d(x) &=& dx&\text{The variable of differentiation}\\ 3&d(|x|) &=& \frac{x}{|x|}~dx&\text{Absolute value function}\\ 4&d(x^c) &=& cx^{c-1}~dx&\text{Power functions}\\ 5&d(\sqrt x) &=& \frac{dx}{2\sqrt x}&\text{Square root function}\\ 6&d(c^x) &=& c^{x}\ln(c)~dx&\text{Exponential functions}\\ 7&d(e^x) &=& e^x~dx&\text{Exponential function}\\ 8&d(\log_c(x)) &=& \frac{dx}{x\ln(c)}&\text{Logarithmic functions}\\ 9&d(\ln(x)) &=& \frac{dx}{x}&\text{Natural logarithmic functions}\\ \hline 10&d(cu) &=& c~du&\text{Constant times a function}\\ 11&d(u+v) &=& du+dv&\text{Sum of functions}\\ 12&d(u\cdot v) &=& v~du+u~dv&\text{Product of functions}\\ 13&d\left(\frac uv\right) &=& \frac{v~du-u~dv}{v^2}&\text{Quotient of functions}\\ 14&d(u\circ v) &=& \frac{du}{dv}~dv&\text{Composition of functions}\\ \end{array} \]

OpenStudy (anonymous):

I just a little explanation of Why and How

OpenStudy (anonymous):

Oh my.

OpenStudy (anonymous):

Start with number \(13\), quotient rule \[ y=\frac{(3x-7)^6}{(2x+1)^3} = \frac uv \]

OpenStudy (anonymous):

We need the derivative of \(u\), in the numerator: \[ u = (3x-7)^6 \]

OpenStudy (anonymous):

6(3x-7)^5

OpenStudy (anonymous):

close, but not quite

OpenStudy (anonymous):

We'll use the composition rule. \[ w=3x-7,\quad u = w^6 \]

OpenStudy (anonymous):

So we think of \(u\) as a function of \(w\). We differentiate with respect to \(w\):\[ du =6w^5 ~dw \]

OpenStudy (anonymous):

Then we find \(dw\): \[ dw = \frac{dw}{dx}~dx = 3~dx \]

OpenStudy (anonymous):

We can plug our \(x\) parts back in: \[ du = 6w^5~dw = 6(3x-7)^5(3~dx) = 18(3x-7)^5~dx \]

OpenStudy (anonymous):

So, at least now we can say: \[ \frac{du}{dx} = 18(3x-7)^5 \]

OpenStudy (anonymous):

Okay. Give me a minute to get this down. This is a little more complex then I was taught.

OpenStudy (anonymous):

We need to find \(dv/dx\), which will be very similar.

OpenStudy (anonymous):

\[ v=(2x+1)^3 \]

OpenStudy (anonymous):

So it would be 6(2x+)^3?

OpenStudy (anonymous):

(2x+1) sorry

OpenStudy (anonymous):

Well, the \(3\) will go to \(2\)

OpenStudy (anonymous):

shoot... that was a little mistake, that's what I meant.

OpenStudy (anonymous):

Okay, assuming that you have the derivatives of \(u\) and \(v\)...\[ dy = \frac{v~du-u~dv}{v^2} \]

OpenStudy (anonymous):

\[\frac{ (2x+1) 18(3x-7)^{5} - (3x - 7) 6(2x+1)^{2} }{ (2x+1)^6 }\]

OpenStudy (anonymous):

close, but your powers are off

OpenStudy (anonymous):

Yes. I forgot to add the powers to my du. (2x+1)^3 and dv (3x -7)^6